# Length independent generalization bounds for deep SSM architectures with stability constraints

### v1

@@ Abstract
Many state-of-the-art models trained on long-range sequences, for example S4, S5 or LRU, are made of sequential blocks combining State-Space Models (SSMs) with neural networks. In this paper we provide a PAC bound that holds for these kind of architectures with stable SSM blocks and does not depend on the length of the input sequence. Imposing stability of the SSM blocks is a standard practice in the literature, and it is known to help performance. Our results provide a theoretical justification for the use of stable SSM blocks as the proposed PAC bound decreases as the degree of stability of the SSM blocks increases.
Keywords: generalization bounds, state-space models, stability, Rademacher complexity
@@ 1 Introduction
The demand for learning powerful representation of time series have been present for decades. The problem of modeling long-range sequences, i.e. sequences with large number of time-steps, is an especially challenging task of the field. Recurrent Neural Networks (RNNs) have always had a hard time dealing with long time series, even advanced architectures such as LSTM or GRU networks did not make a breakthrough. Transformer based architectures managed to imrpove the results in some cases, however still performed poorly on more difficult tasks.
@@
Recently, several novel architectures have been published that are outperforming previous models by a significant margin. The common basis of these models are the so-called Structured State-Space Models (SSMs), which are basically dynamical systems of either continuous or discrete time. Some of the well-performing architectures are the S4 and S5 models and their variants, or another linear system based model called LRU. These models all contain some form of SSMs, either linear or nonlinear, and combine them with some sort of nonlinear transformation from a single nonlinearity to deep neural networks. These blocks of SSMs and neural networks then may be concatenated resulting in a deep structure.
These trending architectures are usually validated empirically on commonly accepted datasets such as the Long Range Arena benchmark. On the other hand, the theoretical foundations of deep, SSM based structured are yet to be discovered. One key point of these models is that they are equipped with some form of stability constraints. For example, the LRU model contains linear SSMs with a certain exponential parametrization of its coefficient matrices enforcing stable behavior of the system. This motivates the question:
What is the role of stability in the success of deep SSM architectures for long-range sequences?
@@
In this paper, we focus on this question and provide a theoretical framework to analyze the model’s generalization behavior in a rigorous manner. We show that stability of deep SSM architectures has an influence on their Rademacher complexity, resulting in a generalization bound that does not depend on the length of the input sequence. In more details:
  • We prove that the Rademacher complexity of a set of SSMs can be upper bounded by the maximal system norm of the considered systems. We consider two types of system norms: $H_{2}$ and $\ell_{1}$ norm. These norms defined for dynamical systems are well-known in control theory, along with their connection to quadratic stability. This result clearly suggests that enforcing stability at least up to a certain level is far beyond being a practical necessity and, in fact, a significant aspect of understanding the essence of these architectures. Consequently, we consider this point as the main takeaway of our paper.
  • Furthermore, we establish an upper bound on the Rademacher complexity of multilayer models by introducing the property of Rademacher Contraction for various components of deep SSM structures, containing multiple separate SSMs combined with nonlinear transformations in between, which is applicable for the popular architectures mentioned above.
  • Using the previous results we establish a PAC-bound on the generalization error of deep SSMs. The resulting bound does not depend on the length of the input sequences, and it only implicitly depends on the depth of the model.
Most of the state-of-the art deep SSM architectures for long range sequences are covered by our results. Moreover, all our our definitions, theorems and proofs hold for both continuous and discrete time systems, as well as for both classification and regression tasks. For the sake of compactness, we stick to the discrete time representations.
Outline of the paper. In Section 2 we present the related literature, then we set notations, definitions and our assumptions related to deep SSMs in Section 3. We propose our main result in Section 4. The majority of the proofs and some additional details are shown in the Appendix.
2 Related work.
@@
There are several PAC bounds for either discrete or continuous-time RNNs in Koiran and Sontag (1998); Sontag (1998); Hanson et al. (2021) by using VC dimension usually through covering numbers. PAC bounds for RNNs based on Rademacher complexity were presented in Wei and Ma (2019); Akpinar et al. (2020); Joukovsky et al. (2021); Chen et al. (2020), while in Zhang et al. (2018) the authors developed PAC-Bayesian bounds. Bounds for general RNNs are related to SSMs as Linear Time-Invariant (LTI) dynamical systems are essential elements almost for all SSMs and they are a special class of RNNs. However all of the previously mentioned PAC bounds depend on the number of time steps and the error bounds tend to infinity with the integration time in continuous case or with the number of time steps in discrete case. That is, these error bounds are not meaningful for prediction errors on long-range sequences.
PAC bounds for both linear and nonlinear dynamical systems are either consider problems not applicable to LTI systems (therefore not applicable for SSMs considered in this paper, see Alquier and Wintenberger (2012); Alquier et al. (2013); Shalaeva et al. (2020); Haussmann et al. (2021); Marion (2023); Haddouche and Guedj (2022)), or not length independent (Hanson and Raginsky (2024)). For example PAC bounds were developed for discrete-time autoregressive models in Alquier and Wintenberger (2012); Alquier et al. (2013); Shalaeva et al. (2020), for stochastic differential equations in Haussmann et al. (2021), for neural ordinary differential equations in Marion (2023), or for online learning in Haddouche and Guedj (2022). In Hanson and Raginsky (2024) the authors propose a PAC bound based on Rademacher complexity for input-affine non-linear systems however their bound is still exponential in the length of the integration interval.
The generalization bound for single vanilla RNNs in (Chen et al., 2020, Theorem 2) is an upper bound of the $H_{1}$ norm proposed in this paper, see e.g. Chellaboina et al. (1999), thus our results based on the $H_{2}$ norm is tighter. Their bound is consistent with our in case of a single linear RNN block. Additionally, they assume bounded matrix norms and their bound depends on the dimension of the representation. In comparison our bound does not depend on the dimension of the state space and we assume only that the systems have bounded $H_{2}$ norm, even in case of stacked RNNs.
In Trauger and Tewari (2024) the authors propose a sequence length independent Rademacher complexity bound for a single layer transformer architecture. For multi layer transformers they improve slightly the result in Edelman et al. (2022) however the bound grows logarithmically with the sequence length.
There are some relevant results outside of the field of RNNs or dynamical systems dealing with the depth of deep neural networks in generalization bounds Bartlett et al. (2017); Liang et al. (2019); Golowich et al. (2018); Truong (2022b). In Golowich et al. (2018) the authors derived a depth independent bound under the condition of bounded Schatten p-norm and a bound with polynomial dependence on depth for Rademacher complexity for DNNs with ReLU activations by applying contraction. Similar approaches utilized spectral norm Bartlett et al. (2017) or Fisher-Rao norm Liang et al. (2019) to establish bounds with non exponential dependence on depth. In a recent paper Truong (2022b) the author extends the bound in Golowich et al. (2018) for non ReLU activations and show that the new, non-vacuous bound is depth independent.
3 Problem setup
In this section, we formally define deep SSM models in way that popular architectures from the literature are contained, the generalization error and the concept of PAC bounds. As our results stand for both continuous-time (CT) and discrete-time (DT) SSMs, in some cases we introduce parallel definitions corresponding to each case.
Notation. We denote scalars with lowercase characters, vectors with lowercase bold characters and matrices with uppercase characters. For a matrix $A$ let $A_{i}$ denote its $i$th row. The symbol $\odot$ denotes the elementwise product. We use $[n]$ to denote the set $\{1,2,\ldots,n\}$ for $n\in\mathbb{N}$.
For vector valued time dependent functions originating from a finite set, the notation $\mathbf{u}_{i}^{(j)}(t)$ refers to the $j$th coordinate of the $i$th function at time $t$. Similarly, in the discrete time case, the notation $\mathbf{u}_{i}^{(j)}[k]$ refers to the $j$th component of the $i$th time series at time step $k$.
$\Sigma$ denotes a dynamical system specified in the context. The constant $n_{\text{in}}$ refers to the dimension of the input sequence, $T$ refers to its length in time, while $n_{\text{out}}$ is the dimension of the output (not necessarily a sequence).
@@ 3 Problem setup
Denote by $\ell^{2,2}_{T}(\mathbb{R}^{n})$ and $\ell^{\infty,\infty}_{T}(\mathbb{R}^{n})$ the Banach spaces generated by the all finite sequences over $\mathbb{R}^{n}$ of length $T$ with the norm $\left\lVert\mathbf{u}\right\rVert^{2}_{\ell^{2,2}_{T}(\mathbb{R}^{n})}=\sum_{k=0}^{T-1}\left\lVert\mathbf{u}[k]\right\rVert_{2}^{2}$ and $\left\lVert\mathbf{u}\right\rVert_{\ell^{\infty,\infty}_{T}(\mathbb{R}^{n})}=\sup\limits_{k=0,\ldots,T-1}\left\lVert\mathbf{u}[k]\right\rVert_{\infty}$ respectively. More generally, denote by $\ell^{2,2}(\mathbb{R}^{n})$ and $\ell^{\infty,\infty}(\mathbb{R}^{n})$ the Banach spaces generated by the all inifinite sequences over $\mathbb{R}^{n}$ such that the quantities $\left\lVert\mathbf{u}\right\rVert^{2}_{\ell^{2,2}(\mathbb{R}^{n})}=\sum_{k=0}^{\infty}\left\lVert\mathbf{u}[k]\right\rVert_{2}^{2}$ and $\left\lVert\mathbf{u}\right\rVert_{\ell^{\infty,\infty}(\mathbb{R}^{n})}=\sup\limits_{k}\left\lVert\mathbf{u}[k]\right\rVert_{\infty}$ are well defined and finite. If $\mathbf{u}=\mathbf{u}[0]\ldots,\mathbf{u}[T-1]$ is a finite sequence of length $T$, then we can interpret it as an infinite sequence $\mathbf{u}=\mathbf{u}[0]\ldots,\mathbf{u}[T-1],0,0,\ldots$; elements of which are zero after the $T$th element. For a Banach space $\mathcal{X}$, $B_{\mathcal{X}}(r)=\{x\in\mathcal{X}\mid\left\lVert x\right\rVert_{\mathcal{X}}\leq r\}$ denotes the ball of radius $r>0$ centered in zero.
We use the symbols $E_{(u,y)\sim\mathcal{D}}$, $P_{(u,y)\sim\mathcal{D}}$, $E_{\mathbf{S}\sim\mathcal{D}^{N}}$ and $P_{\mathbf{S}\sim\mathcal{D}^{N}}$ to denote expectations and probabilities w.r.t. a probability measure $\mathcal{D}$ and its $N$-fold product $\mathcal{D}^{N}$ respectively (see Section 3.2). The notation $\mathbf{S}\sim\mathcal{D}^{N}$ tacitly assumes that $\mathbf{S}\in(\mathcal{U}\times\mathcal{Y})^{N}$, i.e. $\mathbf{S}$ is made of $N$ triplets of input and output trajectories.
3.1 Deep SSMs
@@
A State-Space Model (SSM) is a discrete-time linear dynamical system of the form
$\Sigma\begin{cases}&\mathbf{x}[k]=A\mathbf{x}[k-1]+B\mathbf{u}[k],\leavevmode\nobreak\ \mathbf{x}[0]=0\\ &\mathbf{y}[k]=C\mathbf{x}[k]+D\mathbf{u}[k]\end{cases}$ (1)
@@ A State-Space Model (SSM) is a discrete-time linear dynamical system of the form
where $A\in\mathbb{R}^{n_{x}\times n_{x}},B\in\mathbb{R}^{n_{x}\times n_{u}},C\in\mathbb{R}^{n_{y}\times n_{x}}$ and $D\in\mathbb{R}^{n_{y}\times n_{u}}$ are matrices, is some function of $\mathbf{u}[k],\mathbf{x}[k]$, and $k=1,2,\ldots,T$, where $T$ is the number of time steps. We consider the value of $T$ to be fixed. The reason behind this is rather technical, namely to conviniently handle the Pooling layer (see Definition 6). We emphasize that the generalization bound in Theorem 18 is independent of $T$. For regression tasks the model does not contain any pooling layer, thus this restriction is not needed.
Input-output maps of SSMs. The SSM (1) can be run for finite or infinite number of time steps, and hence it realizes a sequence to sequence transformation. SSM induce as sequence to sequence transformations: they map any input sequence $\mathbf{u}$ of any length (including infinite length) to the unique output sequence $\mathbf{y}$ of the same length, i.e. an SSM $\Sigma$ induces the input-output map of $\Sigma$, denoted by $\mathcal{S}_{\Sigma}$. Moreover, $\mathcal{S}_{\Sigma}$ is causal, $\mathbf{y}[t]$ depends only on the first $t+1$ inputs. The outputs of SSMs are linear combinations of past inputs, in fact, input-output maps can be expressed by a convolution $\mathbf{y}[k]=\mathcal{S}_{\Sigma}(\mathbf{u})[k]=\sum_{j=0}^{k}H_{j}\mathbf{u}(k-j)$, where $H_{0}=D$ and $H_{j}=CA^{j-1}B$, $j>0$. That is, SSMs mix inputs along the time axis, while preserving causality
Stability and input-output maps of SSMs as operators on $\ell^{p,p}$, $p=\infty,2$. For SSMs to be robust transformations of sequences to sequences, one needs stability. Stability is one, if not the most important concept used in dynamical systems and control theory Antoulas (2005). Intuitively, the solutions of a stable system are continuous in the initial state and input. In particular, for stable systems, a small perturbation in past inputs will not result in an increasing error in future outputs.
Definition 1 (Antoulas (2005))

### v2

@@ Abstract
Many state-of-the-art models trained on long-range sequences, for example S4, S5 or LRU, are made of sequential blocks combining State-Space Models (SSMs) with neural networks. In this paper we provide a PAC bound that holds for these kind of architectures with stable SSM blocks and the bound does not depend on the length of the input sequence. Imposing stability of the SSM blocks is a standard practice in the literature, and it is known to help performance. Our results provide a theoretical justification for the use of stable SSM blocks as the proposed PAC bound decreases as the degree of stability of the SSM blocks increases.
@@ 1 Introduction
The problem of modeling long-range sequences, i.e. sequences with large number of time-steps, is an especially challenging task of the field. Recently, several novel architectures (etc. S4 Gu et al. (2021), S4D Gu et al. (2022), S5 Smith et al. (2022), LRU Orvieto et al. (2023)) have been published that are outperforming previous models by a significant margin. The common basis of these models are, combined with some nonlinearity, the so-called Structured State-Space Models (SSMs), which are basically dynamical systems of either continuous or discrete time. One key point of these models is that they are equipped with some form of stability constraints. This motivates the question:What is the role of stability in the success of deep SSM architectures for long-range sequences?
@@
Contribution. In this paper, we focus on this question and provide a theoretical framework to analyze the model’s generalization behavior in a rigorous manner by showing, to our knowledge, the first generalization bound for deep SSMs. We show that stability of deep SSM architectures has an influence on their Rademacher complexity, resulting in a generalization bound that does not depend on the length of the input sequence.
@@
Related work. Bounds for general RNNs are related to SSMs as Linear Time-Invariant (LTI) dynamical systems are essential elements for almost all SSMs and they are a special class of RNNs. There are several PAC bounds for either discrete or continuous-time RNNs in Koiran and Sontag (1998); Sontag (1998); Hanson et al. (2021) by using VC dimension usually through covering numbers. PAC bounds for RNNs based on Rademacher complexity were presented in Wei and Ma (2019); Akpinar et al. (2020); Joukovsky et al. (2021); Chen et al. (2020), while in Zhang et al. (2018) the authors developed PAC-Bayesian bounds. As all of these results tend to infinity with the integration time (number of time steps) they are not meaningful in case of long-range sequences. In Hanson and Raginsky (2024) the authors propose a PAC bound based on Rademacher complexity for input-affine non-linear systems, however their bound is still exponential in the length of the integration interval. The generalization bound for single vanilla RNNs in (Chen et al., 2020, Theorem 2) is an upper bound of the $H_{1}$ norm proposed in this paper, see e.g. Chellaboina et al. (1999), thus our results based on the $H_{2}$ norm is even tighter. In Golowich et al. (2018) the authors derived a depth independent bound under the condition of bounded Schatten p-norm and a bound with polynomial dependence on depth for Rademacher complexity for DNNs with ReLU activations by applying contraction. In a recent paper Truong (2022b) the author extends this bound for non ReLU activations and show that the new, non-vacuous bound is depth independent. Lastly, we mention that in Trauger and Tewari (2024) the authors propose a sequence length independent Rademacher complexity bound for a single layer transformer architecture. For multi layer transformers they improve slightly the result in Edelman et al. (2022) however the bound grows logarithmically with the sequence length.
2 Preliminaries
@@
$\Sigma$ denotes a dynamical system specified in the context. The constant $n_{\text{in}}$ refers to the dimension of the input sequence, $T$ refers to its length in time, while $n_{\text{out}}$ is the dimension of the output (not necessarily a sequence). Denote by $\ell^{2,2}_{T}(\mathbb{R}^{n})$ and $\ell^{\infty,\infty}_{T}(\mathbb{R}^{n})$ the Banach spaces generated by the all finite sequences over $\mathbb{R}^{n}$ of length $T$ with the norm $\left\lVert\mathbf{u}\right\rVert^{2}_{\ell^{2,2}_{T}(\mathbb{R}^{n})}=\sum_{k=0}^{T-1}\left\lVert\mathbf{u}[k]\right\rVert_{2}^{2}$ and $\left\lVert\mathbf{u}\right\rVert_{\ell^{\infty,\infty}_{T}(\mathbb{R}^{n})}=\sup\limits_{k=0,\ldots,T-1}\left\lVert\mathbf{u}[k]\right\rVert_{\infty}$ respectively. For a Banach space $\mathcal{X}$, $B_{\mathcal{X}}(r)=\{x\in\mathcal{X}\mid\left\lVert x\right\rVert_{\mathcal{X}}\leq r\}$ denotes the ball of radius $r>0$ centered in zero.
Generalization gap. We consider the usual supervised learning framework for sequential input data. The considered models, parametrized by $\theta$, are of the form $f_{\theta}:\ell_{T}^{2,2}(\mathbb{R}^{n_{\text{in}}})\to\mathbb{R}^{n_{\text{out}}}$. In this paper, we are agnostic regarding the origin of $\theta$. A dataset is an i.i.d sample of the form $S=\{(\mathbf{u}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ from some distribution $\mathcal{D}$ over $\ell_{T}^{2,2}(\mathbb{R}^{n_{\text{in}}})\times\mathbb{R}$. An elementwise loss function is of the form $\ell:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$. Let $\mathcal{L}_{emp}^{S}(f)=\frac{1}{N}\sum_{i=1}^{N}\ell(f(\mathbf{u}_{i}),y_{i})$ denote the empirical loss of a model $f$ w.r.t a dataset $S$. We denote the true error by $\mathcal{L}(f)=\mathbf{E}_{(\mathbf{u},y)\sim\mathcal{D}}[\ell(f(\mathbf{u}),y)]$. The generalization error or gap of a model $f$ is defined as $|\mathcal{L}_{emp}^{S}(f)-\mathcal{L}(f)|$.
@@
SSMs. A State-Space Model (SSM) is a discrete-time linear dynamical system of the form
$\Sigma\begin{cases}&\mathbf{x}[k]=A\mathbf{x}[k-1]+B\mathbf{u}[k],~{}\mathbf{x}[0]=0\\ &\mathbf{y}[k]=C\mathbf{x}[k]+D\mathbf{u}[k]\end{cases}$ (1)
@@
where $A\in\mathbb{R}^{n_{x}\times n_{x}},B\in\mathbb{R}^{n_{x}\times n_{u}},C\in\mathbb{R}^{n_{y}\times n_{x}}$ and $D\in\mathbb{R}^{n_{y}\times n_{u}}$ are matrices, $\mathbf{u}[k],\mathbf{x}[k]$ and $\mathbf{y}[k]$ are the input, the state and the output signals respectively for $k=1,2,\ldots,T$, where $T$ is the number of time steps. We consider the value of $T$ to be fixed to handle pooling. We emphasize that the generalization bound in Theorem 3.3 is independent of $T$.
Stability. We call the SSM (1) stable, if the matrix $A$ is Schur, i.e., the moduli of all its eigenvalues are smaller than $1$. Intuitively, stable SSMs are robust to perturbations, i.e., their state and outputs are continuous in the initial state and input, see for instance Antoulas (2005) for a more detailed discussion. In particular, a sufficient (but not necessary) condition for stability is that $A$ is a contraction, i.e. $\|A\|_{2}<1$. For more details, see Appendix A.
Input-output maps of SSMs as operators on $\ell^{p,p}_{T}$,$p=\infty,2$. An SSM (1) induces an input-output map, which maps every input sequence $\mathbf{u}[0],\ldots,\mathbf{u}[T-1]$ to output sequences $\mathbf{y}[0],\ldots,\mathbf{y}[T-1]$, which can be interpreted as a linear operator $\mathcal{S}_{\Sigma,T}$ from $\ell^{p,p}_{T}(\mathbb{R}^{n_{u}})\to\ell^{r,r}_{T}(\mathbb{R}^{n_{y}})$, for any choice $p,r\in\{\infty,2\}$. In particular, $\mathcal{S}_{\Sigma,T}$ has a well-defined induced norm as a linear operator, defined in the usual way, $\left\lVert\mathcal{S}_{\Sigma,T}\right\rVert_{p,r}=\sup_{\mathbf{u}\in\ell^{p,p}_{T}(\mathbb{R}^{n_{u}})}\frac{\left\lVert\mathcal{S}_{\Sigma,T}(\mathbf{u})\right\rVert_{\ell^{r,r}_{T}(\mathbb{R}^{n_{y}})}}{\left\lVert\mathbf{u})\right\rVert_{\ell^{r,r}_{T}(\mathbb{R}^{n_{y}})}}$. If $\Sigma$ is internally stable, then it is a standard result in control theory that the norms $\{\left\lVert\mathcal{S}_{\Sigma,T}\right\rVert\}_{T=1}^{\infty}$ are bounded, i.e., $\left\lVert\Sigma\right\rVert_{p,r}=\sup_{T>0}\left\lVert\mathcal{S}_{\Sigma,T}\right\rVert_{p,r}$ exists and it is finite, see Antoulas (2005). Moreover, $\left\lVert\Sigma\right\rVert_{p,r}$ can be viewed as the norm of the extension of the input-output operators $\mathcal{S}_{\Sigma,T}$ to the Banach space generated by $\{\ell^{p,p}_{T}(\mathbb{R}^{n_{u}}\}_{T=1}^{\infty}$ Antoulas (2005). In this paper, we will use the norms $\left\lVert\Sigma\right\rVert_{2,\infty}$ and $\left\lVert\Sigma\right\rVert_{\infty,\infty}$ to upper bound the Rademacher complexity. For more details about the norms see Appendix A.
Relationship with continuous-time models. SSMs are often derived by discretizing a continuous-time linear differential equation in time (e.g. Gu and Dao (2023) and references therein). If the discretization step $\Delta$ is a fixed constant, then we obtain time-invariant linear system of the form (1). Now let us define a single discrete-time SSM block.
Definition DT-SSM. A DT-SSM block (or simply SSM block) is a function $f^{\text{DTB}}:\ell^{p,q}_{T}(\mathbb{R}^{n_{u}})\to\ell^{r,s}_{T}(\mathbb{R}^{n_{u}})$ that is composed of a stable SSM followed by a nonlinear transformation that is constant in time. That is, $f^{\text{DTB}}(\mathbf{u})[k]=g(\mathcal{S}_{\Sigma,T}(\mathbf{u})[k])+\alpha\mathbf{u}[k]$ for some $\alpha\in[0,1]$ and $g:\mathbb{R}^{n_{u}}\to\mathbb{R}^{n_{u}}$ for all $k\in\left[T\right]$.
We incorporate $\alpha$ so that the definition covers residual connections. A deep SSM model consists of SSM blocks along with an encoder, and a decoder transformation preceded by a time-pooling layer. We present an overview of the various architectures found in the literature in Table 1, Appendix B. Now we may define deep SSMs. For precise details about the particular elements see Appendix B.
Definition DT deep SSM. A discrete time deep SSM model for classification is a function $f:\ell^{p,q}_{T}(\mathbb{R}^{n_{\text{in}}})\to\mathbb{R}^{n_{\text{out}}}$ of the form $f=f^{\text{Dec}}\circ f^{\text{Pool}}\circ f^{\text{B}_{L}}\circ\ldots\circ f^{\text{B}_{1}}\circ f^{\text{Enc}}$, where $\circ$ denotes composition of functions. The functions $f^{\text{Enc}}$ and $f^{\text{Dec}}$ are linear transformations which are constant in time, while $f^{\text{B}_{i}}$ is a DT-SSM block for all $i$. By pooling we mean the operation $f^{\text{Pool}}(\mathbf{u})=\frac{1}{T}\sum_{k=1}^{T}\mathbf{u}[k]$, an average pooling over the time axis.
3 Rademacher contraction of deep SSMs
Before we state our main theorem we introduce a property of functions, referred to as Rademacher Contraction, that is universal enough to include functions represented by both deep SSMs and neural networks.
Definition 3.1 ($(\mu,c)$-Rademacher Contraction).
Let $X_{1}$ and $X_{2}$ be subsets of Banach spaces $\mathcal{X}_{1},\mathcal{X}_{2}$, with norms $\|\cdot\|_{\mathcal{X}_{1}}$ and $\|\cdot\|_{\mathcal{X}_{2}}$, and let $\mu\geq 0$ and $c\geq 0$. A set of functions $\Phi=\{\varphi:X_{1}\to X_{2}\}$ is said to be $(\mu,c)$-Rademacher Contraction, or $(\mu,c)$-RC in short., if for all $n\in\mathbb{N}^{+}$ and $Z\subseteq X_{1}^{n}$ we have
The RC property can be used to upper bound the Rademacher complexity of the hypothesis class. In fact, special cases of Definition 3.1 were used in the literature to this effect Golowich et al. (2018); Truong (2022a); Trauger and Tewari (2024) for deep neural networks and transformers respectively.
All blocks of deep SSMs have the RC property, for the proofs see Appendix D.1. In particular, SSM components of SSM blocks are $(\mu,0)$-RC, and the constants $\mu$ depends on the control-theoretic system norm of the SSM component. Moreover, the RC property is preserved under composition of layers.
Lemma 3.2 (Composition lemma).
The proof is in Appendix D.2. Lemma (3.2) allows us to establish the RC property for any deep structure, in particular, for deep SSMs. Next we describe our assumptions for deep SSMs.
Let $\mathcal{F}$ be a set of deep SSM models, namely let $f\in\mathcal{F}$ has the form $f=f^{\text{Dec}}\circ f^{\text{Pool}}\circ f^{\text{B}_{L}}\circ\ldots\circ f^{\text{B}_{1}}\circ f^{\text{Enc}}$. Let us assume that (1) there exist constants $K_{\mathbf{u}}$ and $K_{y}$ such that with probability one w.r.t. the data distribution $\mathcal{D}$, for any input-label pair $(\mathbf{u},y)$, the $\ell^{2,2}_{T}$ norm of $\mathbf{u}$ and the absolute value of $y$ are bounded from above $K_{\mathbf{u}}$ and $K_{y}$ respectively, (2) the elementwise loss is $L_{l}$-Lipschitz continuous, (3) the SSM component of $\Sigma_{i}$ of each SSM block in the parametrization is stable and its norm $\|\Sigma_{i}\|_{p,q}$ is bounded by a constant $K_{p,q}$, $p,q=2,\infty$, (4) the norms of the weights of the encoder $f^{\text{Enc}}$ and decoder $f^{\text{Dec}}$ are bounded by $K_{\text{Enc}},K_{\text{Dec}}$ respectively, (5) the non-linear component $g_{i}$ of the $i$th SSM block $g_{i}$ is $(\mu_{g_{i}},c_{g_{i}})$-RC. Internal stability of the SSM components is a standard assumption in the literature, and it is shown in Appendix D that the commonly used non-linear components are RC. We present the assumptions formally in Appendix C. Now we state our main theorem.
Theorem 3.3 (Informal theorem).
The following PAC inequality holds
$\mathbb{P}_{S\sim\mathcal{D}^{N}}\left[\forall f\in\mathcal{F}\quad\mathcal{L}(f)-\mathcal{L}_{emp}^{S}(f)\leq\frac{\mu K_{\mathbf{u}}L_{l}+cL_{l}}{\sqrt{N}}+K_{l}\sqrt{\frac{2\log(4/\delta)}{N}}\right]>1-\delta,$
where the constants $\mu$ and $c$ depend on the hypothesis class $\mathcal{F}$ and they satisfy
$\mu\leq K_{\text{Enc}}K_{\text{Dec}}\prod\limits_{i=1}^{L}\left(\mu_{g_{i}}K+\alpha_{i}\right),$ $c\leq K_{\text{Dec}}\sum\limits_{j=1}^{L}\left[\prod\limits_{i=j+1}^{L}\left(\mu_{g_{i}}K+\alpha_{i}\right)\right]c_{g_{j}}$
and the constant $K_{l}>0$ such that $|l(\cdot,\cdot)|\leq K_{l}$, while $K=\max\{K_{2,\infty},K_{\infty,\infty}\}$.
The formal counterpart of Theorem 3.3 and its proof can be found in Appendix E. Notice that the bound does not depend on $T$. While the bound grows with the depth of the deep SSM, its growth can be controlled by choosing the state-space blocks with a small system norm. In turn, the system norm of the state-space models depends not only on the number and magnitude of its parameters, but on the degree of stability, i.e., systems with large number of weights with possibly large parameter norms can still have a small system norm. In practice, for popular deep SSM architectures, e.g. S4, S4D, S5 or LRU, the stability conditions are naturally met due their constrained parametrizations. This indicates that stability induced norms are crucial for deep SSMs.
4 Ackowledgements
This research was supported by the European Union project RRF-2.3.1-21-2022-00004 within the framework of the Artificial Intelligence National Laboratory and by the C.N.R.S. E.A.I. project "Stabilité des algorithmes d’apprentissage pour les réseaux de neurones profonds et récurrents en utilisant la géométrie et la théorie du contrôle via la compréhension du rôle de la surparamétrisation".
Appendix A Stability of SSMs
We recite the definition of internally stable dynamical systems of the form eq. 1.
Definition A.1 (Antoulas (2005)).